OpenStudy (anonymous):

Evaluate the given expression a = -4, b = 8, c = 2, and d = -3

OpenStudy (anonymous):

The expression to be evaluated is missing.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

|ad-c|/|b-3|

OpenStudy (anonymous):

@RolyPoly

OpenStudy (anonymous):

please quickly running out of time

OpenStudy (anonymous):

|ad-c|/|b-3| First, let us consider the numerator. |ad-c| Substitute a = -4, c = 2, and d = -3 into the expression |ad-c| What is (-4)(-3)-(2) ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Now consider |b-3| Substitute b=8 into b-3 What is 8-3?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Right, so |ad-c|/|b-3| = 10/5 = ...?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Yes :)